a.     Alemu purchased office supplies worth $5,000 from various suppliers agreeing to pay the sum within two weeks (Purchase invoice #1). Two-fifth of the supplies will be used for personal purposes while the remaining is for use by AH. Three-fifth of the liability is arranged to be settled from business cash sources.  



1
Expert's answer
2021-10-18T12:05:37-0400

Total office supplies = $5,000

Drawings (personal use) =2/5*5,000=$2,000

AH supplies =3/5*5,000=$3,000

Arranged for settlement on business cash

Dr. Sundry creditors 3,000

Cr. Cash a/c                          3,000